OpenStudy (anonymous):

Find the slope of the tangent line to the curve y=logBase5(3x-1) at point x=3

OpenStudy (anonymous):

I 'll try to help but i see satelite 73 here, let him helps you

OpenStudy (anonymous):

Hoa-I just need to know how to start it any help will do(:

OpenStudy (anonymous):

go ahead the idea is to take the derivative and replace \(x\) by 3 what you need to know is that \[\frac{d}{dx}\log_b(x)=\frac{1}{x\ln(b)}\]

OpenStudy (anonymous):

that plus the chain rule should get you the derivative if that is not clear let me know

OpenStudy (anonymous):

Okay let me give it a try, Thanks(:

OpenStudy (anonymous):

don't forget the "x" here is 3x-1 not just x . I don't know whether it's right or wrong. my pro leads me another way when solving that problem. ok. whatever, go to his way first.

OpenStudy (anonymous):

I got it, Thanks to both of you(: